At what velocity will this drop of dextrose hit the floor?




1
Expert's answer
2021-11-19T10:50:34-0500

We can find the velocity from the kinematic equation:


"v=v_0+gt=0+(-9.8\ \dfrac{m}{s^2}\times0.68\ s)=-6.7\ \dfrac{m}{s}."

The sign minus means that the velocity of the drop directed downward.